# OpenReward: Learning to Reward Long-form Agentic Tasks via Reinforcement Learning

### v1

@@ 1 Introduction
As illustrated in Table 1, a key obstacle in training long-form reward models is the lack of reliable pairwise response data from real-world sources to serve as training signals. To address this, we propose a simple yet controllable partial synthesis framework that automatically generates queries and constructs positive–negative pairs. Specifically, we first perform target-aware query generation, prompting a strong LLM (e.g., DeepSeek-V3) to formulate a self-contained query from a sampled document. Then, we obtain preference-ordered responses by prompting the LLM answer the query with and without access to the reference document, thereby yielding a positive and a negative response, respectively. This method enables scalable supervision without costly human annotations.
Table 1: Basic statistics of our benchmark.
As a further step, we explore the utility of OpenRM in scaling LLM alignment at both training and inference stages. At inference time, we show that OpenRM can reliably select the better response among candidates, improving long-form answer quality. At training time, we employ OpenRM as a data selector, where we compare task-solving trajectories sampled from LLMs and filtering with OpenRM, thereby obtaining a higher-quality training dataset. Our experiments demonstrate that LLMs trained with OpenRM-selected data achieve significant performance gains compared with those trained on data filtered by existing RMs, highlighting the potential of tool-augmented reward modeling.
@@ 4 Experiment Setup
Domain Wikipedia Scientific Medical Average Direct reasoning without tool utilization. Deepseek-V3.1 (Liu et al., 2024a) 75.00 46.00 33.00 51.33 GPT-4o (Hurst et al., 2024) 70.00 48.20 44.00 54.07 Gemini-2.5-Pro (Comanici et al., 2025) 72.20 46.60 36.00 51.60 Claude-Opus-4-1-20250805 (Anthropic Team, 2025) 74.60 49.20 51.10 58.30 Skywork-Reward-Gemma-2-27B (Liu et al., 2024b) 45.20 55.40 47.74 49.45 JudgeLRM-7B (Chen et al., 2025a) 50.80 50.60 48.44 49.94 RRM-7B (Guo et al., 2025) 56.90 52.95 53.10 54.32 RM-R1-Qwen2.5-Instruct-7B (Chen et al., 2025b) 55.40 54.80 52.30 54.17 RM-R1 trained on our data 66.00 73.00 65.00 68.00 Reasoning with tool utilization. Deepseek-V3.1 (Liu et al., 2024a) 77.00 48.00 34.00 53.00 GPT-4o (Hurst et al., 2024) 76.40 58.60 53.40 62.80 Gemini-2.5-pro (Comanici et al., 2025) 72.50 54.60 42.40 56.50 Claude-Opus-4-1-20250805 (Anthropic Team, 2025) 75.60 56.10 58.00 63.23 OpenRM-Qwen-2.5-7B 93.00 90.00 91.00 91.33
Table 2: Comparison between OpenRM and baseline models in terms of Accuracy ($\uparrow$) across three in-domain tasks (Wikipedia, Scientific, and Medical). We report results both under direct reasoning without external tool usage and under reasoning with tool utilization. OpenRM consistently outperforms all baselines, achieving the highest average accuracy (91.33), which highlights the effectiveness of combining reasoning with external retrieval.
Methods Backbone Data ($\downarrow$) PandaLM RewardBench Average Average $\Delta$ Out-of-domain Evaluation OpenRM Qwen-2.5-7B-Instruct 27K 79.42 77.66 78.54 - RM-R1 Qwen-2.5-7B-Instruct 72K 72.71↓6.71 68.34↓9.32 70.52 11.37 JudgeLRM-7B Qwen-2.5-7B-Instruct 100K 72.37↓7.05 74.45↓3.21 73.41 $\downarrow$ 6.99 Prometheus-v2.0 Mistral-7B-Instruct 200K 72.80↓6.62 71.55↓6.11 72.17 $\downarrow$8.83 RRM-7B Qwen-2.5-7B-Instruct 420K 77.74↓1.68 78.54↑0.88 78.14 $\downarrow$0.52
@@ 5.1 Overall performance
Figure 2: Illustration of the overlap between RewardBench and the training datasets of different reward models. The embeddings are visualized in 2D, where orange points denote training data and blue points denote RewardBench instances. We observe substantial overlap in JudgeLRM, RRM, and RM-R1, whereas our dataset shows almost no intersection, ensuring a fairer and less biased evaluation.
@@ 5.2 Ablation studies
  1. 1.
    w/ only $\mathcal{R}_{\text{EM}}$, which uses only the judgment accuracy signal $\mathcal{R}_{\text{EM}}\in[0,1]$.
  2. 2.
    w/ $\mathbf{1}_{\mathcal{R}_{\text{EM}}}\cdot\mathcal{R}_{\text{tool}}$, where the weight of $\mathcal{R}_{\text{tool}}$ is increased from $0.5$ to $1$, formulated as $\mathcal{R}_{\text{EM}}+\text{sign}(\mathcal{R}_{\text{EM}})\cdot\mathcal{R}_{\text{tool}}\in[0,2]$;
  3. 3.
    w/ $0.5\cdot\mathcal{R}_{\text{tool}}$: removing the indicator $\mathbf{1}_{\mathcal{R}_{\text{EM}}}$ when supervising tool usage, formulated as $\mathcal{R}_{\text{EM}}+\mathcal{R}_{\text{tool}}\in[-0.5,1.5]$. This decouples tool-use learning from judgment learning, encouraging more explicit tool usage.
First, Second, Third, Since these reward functions operate on different scales, we normalize them to a unified $[0,1]$ range for fair comparison.
@@
Methods Backbone Self-contain Factuality Average Average $\Delta$ Metrics Accuracy ($\uparrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$) DeepSeek-V3.1 (Liu et al., 2024a) - 2.67 2.47 2.57 - OpenRM Qwen-2.5-7B-Instruct 2.73 2.83 2.78 0.21 RM-R1 Chen et al. (2025b) Qwen-2.5-7B-Instruct 2.13 1.73 1.93 -0.64
@@ 6 Utility study
Methods TrutuFulQA ($\uparrow$) MMLU-Pro($\uparrow$) Triviaqa ($\uparrow$) Average ($\uparrow$) Average $\Delta$ ($\uparrow$) Qwen-2.5-3B-Inst. 32.19 38.17 42.33 37.56 - Qwen-2.5-7B-Inst. 39.04 48.55 50.19 45.93 - Training: DPO w/-w/ RM-R1-Qwen-3B-Inst 32.07 38.23 43.69 37.99 0.44 w/-w/ OpenRM-Qwen-3B-Inst 34.03 38.60 44.56 39.03 1.47 w/-w/ RM-R1-Qwen-7B-Inst 41.00 48.81 51.53 47.11 1.18 w/-w/ OpenRM-Qwen-7B-Inst 43.33 50.63 50.12 48.03 2.1
@@
Table 5: Experiment results in terms of accuracy on three out-of-domain benchmarks. We compare Qwen-2.5-3B-Instruct and Qwen-2.5-7B-Instruct trained with either RM-R1 or OpenRM as reward models under direct preference optimization (DPO)
@@
In this section, we demonstrate the practical utility of OpenRM in LLM training, extending its use beyond standalone evaluation. We experiment on the widely used UltraFeedback dataset Cui et al. (2023), which consists of diverse prompts paired with multiple candidate responses. We apply OpenRM to score and filter training data, allowing the model to learn from higher-quality responses and reducing noise in supervision. Additionally, we conduct qualitative case studies to better understand the reasoning capabilities of OpenRM. These studies reveal that OpenRM is more effective at orchestrating tool usage to handle complex tasks. A representative example is shown in Table 6
@@
Dataset: UltraFeedback Question: How have foreign influences played a role in shaping Romania’s cultural identity? Answer A: Foreign influences have played an important role in shaping Romania’s cultural identity. Due to its geographic location, Romania has been exposed to various cultures for centuries, from the neighboring Hungarians and Austrians to further away Ottomans and Russians. As a result, Romanian culture has incorporated aspects from all of these cultures, ranging from cuisine to religious practices to linguistic norms. For example, the traditional Romanian New Year’s Eve dinner includes ntbura, a traditional soup also known as ”Sultan’s soup,” likely due to its Ottoman origins… Answer B: The influence of foreign cultures has played a significant role in shaping Romania’s cultural identity throughout history. Here are some examples: 1. Greek Influence: Ancient Greece had a profound impact on Romanian culture… 2. Ottoman Influence: The Ottoman Empire had a significant influence on Romania’s cultural identity… 3. German Influence: Germany has had a significant cultural impact on Romania… 4. French Influence: France has had a significant cultural impact on Romania… 5. Italian Influence: Italy has had a significant cultural impact on Romania… RM-R1 Judgment: Answer B is better due to its systematic, numbered breakdown of five specific cultural influences (Greek, Ottoman, etc.). This provides a more complete and structured inventory than Answer A’s general narrative, enhancing clarity. This organization is more useful for a user seeking a comprehensive list and easy-to-digest information. <answer><answer>B</answer></answer> ✗ Wrong answer, the explanation for each is very brief and lacks depth. Only focus on the structure, the content is neglected. OPENRM Judgment: Turn 1: Both answers provide information about the role of foreign influences in shaping Romania’s cultural identity. However, Answer A is more detailed, providing concrete examples such as the integration of loanwords into the Romanian language. I will perform a search to see if there are any critical steps missed by either answer, before making a final decision. <search><search>WIKI(foreign influences on Romanian culture)</search></search> Turn 2: The search results returned content related to Slavic influences on Romanian culture, consistent with the description in Answer A. Answer A is more detailed and specific, providing concrete examples such as the Ottoman influence on traditional dishes and the integration of loanwords into the Romanian language. Answer B, while informative, is more structured as a list of examples and lacks the same level of detail and specificity. Therefore, Answer A is more comprehensive and better answers the input question. <answer><answer>A</answer></answer> ✓ Correct answer, A search is conducted to assist in judgment, and it is successfully identified that the content of Answer A is more detailed with more specific examples.
@@
We further apply OpenRM to Direct Preference Optimization (DPO) Rafailov et al. (2023), a widely adopted reinforcement learning method for aligning LLMs with human feedback (i.e., RLHF). This experiments evaluate whether the high-quality responses identified by OpenRM can enhance downstream alignment. In this setting, the previously identified positive and negative responses are used as preference pairs for DPO training. As shown in Table 5, the 3B model trained on data curated by OpenRM achieves 39.03% and the 7B model trained on data curated by OpenRM achieves 48.03% in terms of accuracy on preference prediction, representing a substantial improvement over the vanilla backbone without post-training. Moreover, we find that it surpasses the counterpart trained on data filtered by RM-R1. This result demonstrates that the higher-quality preference data identified by OpenRM can more effectively guide preference-based optimization, leading to stronger alignment. These findings highlight the broader potential of tool-augmented reward modeling: by producing more accurate and reliable preference signals, OpenRM strengthens the effectiveness of downstream alignment algorithms such as DPO, thereby contributing to more robust and human-aligned LLMs.
@@ 7 Conclusion
Despite the promising resutls, our study still has several limitations. First, the effectiveness of OpenRM relies on the availability and reliability of external tools, which may introduce bias or extra latency in evidence retrieval. Second, our current implementation only focuses on text-based evaluation, and has yet to be extended to multimodal settings involving visual or tabular inputs. Looking forward, we plan to extend our framework to more complex scenarios involving a broader set of external tools, thereby enhancing its applicability to diverse tasks. We also aim to adapt OpenRM to multimodal settings, enabling models to seamlessly integrate textual reasoning with visual and tabular evidence. Such extensions would pave the way for building more general and robust evaluation agents capable of aligning large language models with complex real-world requirements.

### v2

@@ 1 Introduction
A key obstacle in training long-form reward models is the lack of reliable pairwise response data from real-world sources to serve as training data, as demonstrated in Table 1. To address this, we propose a simple yet controllable partial synthesis framework that automatically generates queries and constructs positive–negative pairs. Specifically, we first perform target-aware query generation, prompting a strong LLM (e.g., DeepSeek-V3) to formulate a self-contained query from a sampled document. Then, we obtain preference-ordered responses by prompting the LLM answer the query with and without access to the reference document, thereby yielding a positive and a negative response, respectively. This method enables scalable supervision without costly human annotations.
As a further step, we explore the utility of OpenRM in scaling LLM alignment at both training and inference stages. At inference time, we show that OpenRM can select a better response among candidates, improving long-form answer quality. At training time, we employ OpenRM as a data selector, where we compare task-solving trajectories sampled from LLMs and filtering with OpenRM, thereby obtaining a higher-quality training dataset. Our experiments demonstrate that LLMs trained with OpenRM-selected data achieve significant performance gains compared with those trained on data filtered by existing RMs, highlighting the potential of tool-augmented reward modeling.
@@ 4 Experiment Setup
Domain Wikipedia ($\uparrow$) Scientific ($\uparrow$) Medical ($\uparrow$) Average ($\uparrow$) Direct reasoning without tool utilization. Deepseek-V3.1 (Liu et al., 2024a) 75.00 46.00 33.00 51.33 GPT-4o (Hurst et al., 2024) 70.00 48.20 44.00 54.07 Gemini-2.5-Pro (Comanici et al., 2025) 72.20 46.60 36.00 51.60 Claude-Opus-4-1-20250805 (Anthropic Team, 2025) 74.60 49.20 51.10 58.30 Skywork-Reward-Gemma-2-27B (Liu et al., 2024b) 45.20 55.40 47.74 49.45 JudgeLRM-7B (Chen et al., 2025a) 50.80 50.60 48.44 49.94 RRM-7B (Guo et al., 2025) 56.90 52.95 53.10 54.32 RM-R1-Qwen2.5-Instruct-7B (Chen et al., 2025b) 55.40 54.80 52.30 54.17 RM-R1 trained on our data 66.00 73.00 65.00 68.00 Reasoning with tool utilization. Deepseek-V3.1 (Liu et al., 2024a) 77.00 48.00 34.00 53.00 GPT-4o (Hurst et al., 2024) 76.40 58.60 53.40 62.80 Gemini-2.5-pro (Comanici et al., 2025) 72.50 54.60 42.40 56.50 Claude-Opus-4-1-20250805 (Anthropic Team, 2025) 75.60 56.10 58.00 63.23 OpenRM-Qwen-2.5-7B 93.00 90.00 91.00 91.33
Table 2: Comparison between OpenRM and baseline models in terms of Accuracy ($\uparrow$) across three in-domain tasks (Wikipedia, Scientific, and Medical). We report results both under direct reasoning without external tool usage and under reasoning with tool utilization.
@@ 5.1 Overall performance
Methods Backbone Data ($\downarrow$) PandaLM ($\uparrow$) RewardBench ($\uparrow$) Average ($\uparrow$) Average $\Delta$% ($\uparrow$) Out-of-domain Evaluation OpenRM Qwen-2.5-7B-Instruct 27K 79.42 77.66 78.54 - RM-R1 Qwen-2.5-7B-Instruct 72K 72.71↓6.71 68.34↓9.32 70.52 $\downarrow$11.37% JudgeLRM-7B Qwen-2.5-7B-Instruct 100K 72.37↓7.05 74.45↓3.21 73.41 $\downarrow$6.99% Prometheus-v2.0 Mistral-7B-Instruct 200K 72.80↓6.62 71.55↓6.11 72.17 $\downarrow$8.83% RRM-7B Qwen-2.5-7B-Instruct 420K 77.74↓1.68 78.54↑0.88 78.14 $\downarrow$0.52%
@@
Figure 2: Illustration of the overlap between RewardBench and the training datasets of different reward models. The embeddings are visualized in 2D space. The orange points denote training data and blue points denote RewardBench instances. We observe substantial overlap in JudgeLRM, RRM, and RM-R1, whereas our dataset shows almost no intersection, ensuring a fairer and less biased evaluation .
@@ 5.2 Ablation studies
  1. (1)
    only $\mathcal{R}_{\text{EM}}$: we use only the judgment accuracy signal $\mathcal{R}_{\text{EM}}\in[0,1]$.
  2. (2)
    $\mathbf{1}_{\mathcal{R}_{\text{EM}}}\cdot\mathcal{R}_{\text{tool}}$: the weight of $\mathcal{R}_{\text{tool}}$ is increased from $0.5$ to $1$, formulated as $\mathcal{R}_{\text{EM}}+\text{sign}(\mathcal{R}_{\text{EM}})\cdot\mathcal{R}_{\text{tool}}\in[0,2]$;
  3. (3)
    $0.5\cdot\mathcal{R}_{\text{tool}}$: we remove the indicator $\mathbf{1}_{\mathcal{R}_{\text{EM}}}$ when supervising tool usage, formulated as $\mathcal{R}_{\text{EM}}+\mathcal{R}_{\text{tool}}\in[-0.5,1.5]$. This decouples tool-use learning from judgment learning, encouraging more explicit tool usage.
Since these reward functions operate on different scales, we normalize them to a unified $[0,1]$ range for fair comparison.
@@ 5.3 Human evaluation
Methods Backbone Self-contain ($\uparrow$) Factuality ($\uparrow$) Average ($\uparrow$) Average $\Delta$% ($\uparrow$) Metrics Accuracy ($\uparrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$) OpenRM Qwen-2.5-7B-Instruct 2.73 2.83 2.78 - LLM-as-Judge DeepSeek-V3.1 2.67 2.47 2.57 $\downarrow$7.55% RM-R1 Chen et al. (2025b) Qwen-2.5-7B-Instruct 2.13 1.73 1.93 $\downarrow$30.58%
@@ 6 Utility study
Methods TrutuFulQA ($\uparrow$) MMLU-Pro($\uparrow$) Triviaqa ($\uparrow$) Average ($\uparrow$) Average $\Delta$% ($\uparrow$) Out-of-domain Evaluation Qwen-2.5-3B-Instruct 32.19 38.17 42.33 37.56 - w/-w/ RM-R1 32.07 38.23 43.69 37.99 $\uparrow$1.14% w/-w/ OpenRM 34.03 38.60 44.56 39.03 $\uparrow$3.91% Qwen-2.5-7B-Instruct 39.04 48.55 50.19 45.93 - w/-w/ RM-R1 41.00 48.81 51.53 47.11 $\uparrow$2.57% w/-w/OpenRM 43.33 50.63 50.12 48.03 $\uparrow$4.57%
@@
Table 5: Experiment results in terms of accuracy on three out-of-domain benchmarks. We compare Qwen-2.5-3B-Instruct and Qwen-2.5-7B-Instruct trained with either RM-R1 or OpenRM as reward models under direct preference optimization (DPO).
@@
In this section, we demonstrate the practical utility of OpenRM in LLM training, extending its use beyond standalone evaluation. We experiment on the widely used UltraFeedback dataset Cui et al. (2023), which consists of diverse prompts paired with multiple candidate responses.
OpenRM for data selection. We apply OpenRM to score and filter training data, allowing the model to learn from higher-quality responses and reducing noise in supervision. Table 6 shows the selection accuracy on the UltraFeedback dataset, where the accuracy is measured by how often the model selects the response originally labeled better. OpenRM achieves 75.18%, outperforming RM-R1 Chen et al. (2025b) with 68.87%, indicating stronger alignment with human preferences. Additionally, we conduct qualitative case studies to better understand the reasoning capabilities of OpenRM. These studies reveal that OpenRM is more effective at orchestrating tool usage to handle complex tasks. A representative example is shown in Table 7.
Dataset UltraFeedback Metric Accuracy ($\uparrow$) OpenRM 75.18 RM-R1 Chen et al. (2025b) 68.87
Table 6: Selection results accuracy comparing OpenRM (75.18) and RM-R1 Chen et al. (2025b) (68.87) on the UltraFeedback dataset.
@@
Dataset: UltraFeedback Question: How have foreign influences played a role in shaping Romania’s cultural identity? Answer A: Foreign influences have played an important role in shaping Romania’s cultural identity. Due to its geographic location, Romania has been exposed to various cultures for centuries, from the neighboring Hungarians and Austrians to further away Ottomans and Russians. As a result, Romanian culture has incorporated aspects from all of these cultures… For example, the traditional Romanian New Year’s Eve dinner includes ntbura… Answer B: The influence of foreign cultures has played a significant role in shaping Romania’s cultural identity throughout history. Here are some examples: 1. Greek Influence: Ancient Greece had a profound impact on Romanian culture… 2. Ottoman Influence… 3. German Influence… 4. French Influence… 5. Italian Influence… RM-R1 Judgment: Answer B is better due to its systematic, numbered breakdown of five specific cultural influences (Greek, Ottoman, etc.). This provides a more complete and structured inventory than Answer A’s general narrative, enhancing clarity. This organization is more useful for a user seeking a comprehensive list and easy-to-digest information. <answer><answer>B</answer></answer> ✗ Wrong answer: the explanation for each response is brief and lacks in-depth analysis, which only focus on the structure while neglecting content factuality. OpenRM Judgment: Both answers provide information about the role of foreign influences in shaping Romania’s cultural identity. However, Answer A is more detailed, providing concrete examples such as… I will perform a search to see if there are any critical steps missed by either answer, before making a final decision. <<search>>WIKI(foreign influences on Romanian culture)<<search>>The search results returned content related to Slavic influences on Romanian culture, consistent with the description in Answer A. Answer A is more detailed and specific, providing concrete examples such as … Answer B, while informative, is more structured as a list of examples and lacks the same level of detail and specificity. Therefore, Answer A is more comprehensive and better answers the input question. <<answer>>A<<answer>> ✓ Correct answer: a search tool is invoked to assist judgment. The model eventually identifies that the content of Answer A is more detailed than Answer B.
@@
OpenRM for model alignment. We further apply OpenRM to Direct Preference Optimization (DPO) Rafailov et al. (2023), a widely adopted reinforcement learning method for aligning LLMs with human feedback (i.e., RLHF). This experiments evaluate whether the high-quality responses identified by OpenRM can enhance downstream alignment. In this setting, the previously identified positive and negative responses are used as preference pairs for DPO training. As shown in Table 5, the 3B model trained on data curated by OpenRM achieves 39.03% and the 7B model trained on data curated by OpenRM achieves 48.03% in terms of accuracy on preference prediction, representing a substantial improvement over the vanilla backbone without post-training. Moreover, we find that it surpasses the counterpart trained on data filtered by RM-R1. This result demonstrates that the higher-quality preference data identified by OpenRM can more effectively guide preference-based optimization, leading to stronger alignment. These findings highlight the broader potential of tool-augmented reward modeling: by producing more accurate and reliable preference signals, OpenRM strengthens the effectiveness of downstream alignment algorithms such as DPO, thereby contributing to more robust and human-aligned LLMs.
@@ 7 Conclusion
Despite the promising results, our study still has two limitations. First, the effectiveness of OpenRM relies on the availability and reliability of external tools, which may introduce bias or extra latency in evidence retrieval. Second, our current implementation only focuses on text-based evaluation, and has yet to be extended to multimodal settings involving visual or tabular inputs. Looking forward, we plan to extend our framework to more complex scenarios involving a broader set of external tools, thereby enhancing its applicability to diverse tasks. We also aim to adapt OpenRM to multimodal settings, enabling models to seamlessly integrate textual reasoning with visual and tabular evidence. Such extensions would pave the way for building more general and robust evaluation agents capable of aligning large language models with complex real-world requirements.